# Design of Bayesian Clinical Trials with Clustered Data and Multiple Endpoints

### v2

@@ 4 Sample Size Determination Procedure
Bootstrap confidence intervals for the design operating characteristics at a given value of $c$ can similarly be constructed. For each of the $M$ sets of bootstrap samples, the linear approximations obtained using the process in Lines 7 to 8 of Algorithm 1 give rise to a new estimate of the operating characteristics. A bootstrap confidence interval for the operating characteristic can also be calculated using the percentile method. In Section 5, we consider the performance of Algorithm 1 and construct bootstrap confidence intervals for a substantive example motivated by a current clinical trial.
5 Application to the SSTaRLeT Trial
We now apply our proposed approach to the Shorter and Safer Treatment Regimens for Latent Tuberculosis (SSTaRLeT) trial (Menzies, 2024). This trial uses an adaptive platform non-inferiority design where decision criteria are specified with respect to three binary endpoints for the interim analysis. The goal of the trial is to establish non-inferiority of new tuberculosis preventive medications with respect to the reference treatment in terms of safety. Safety is defined using the primary outcome of a grade 3 to 5 adverse event (AE) and two secondary outcomes, completion of the treatment by patients as well as non-tolerability defined as the occurrence of a set of reactionary symptoms. Only the primary outcome is considered at the trial’s final analysis.
We demonstrate the utility of our approach by considering a pairwise comparison between a new treatment and the reference treatment. We emphasize that in the context of an adaptive design, our method proposed in Section 4 is crucial to efficiently estimate the joint sampling distribution of posterior probabilities based on clustered data across all estimands. We shortly describe how to extend our method to account for the adaptive nature of the trial using theory from Hagar et al. (2025). Because our SSD framework is computationally efficient, we can consider complex analysis models to design SSTaRLeT and evaluate a broad set of within-cluster dependence structures, underscoring the practical benefits of our approach.
SSTaRLeT is a cluster-randomized trial where all members of a given household are randomized to the same treatment arm. The trial considers $K=3$ estimands based on the binary endpoints. The first estimand $\delta_{1}(\boldsymbol{\theta})$ is the difference in the marginal probabilities of experiencing an AE between the new ($A=1$) and reference ($A=0$) treatment arms. Estimands $\delta_{2}(\boldsymbol{\theta})$ and $\delta_{3}(\boldsymbol{\theta})$ are respectively the differences between marginal probabilities of non-completion and non-tolerability in the new and reference treatments. For estimands $k=1$ to $3$, the trial subhypotheses are defined as follows:
$H_{0,k}:\delta_{k}(\boldsymbol{\theta})\geq\delta_{U,k}\hskip 5.0pt\text{vs.}\hskip 5.0ptH_{1,k}:\delta_{k}(\boldsymbol{\theta})<\delta_{U,k},$
where $\delta_{U,k}$ is the non-inferiority margin for subhypothesis $k$. The non-inferiority margins for this trial are $\delta_{U,1}=0.04$ and $\delta_{U,2}=\delta_{U,3}=0.1$. Thus, large positive differences between probabilities of experiencing an AE, non-completion, and non-tolerability are associated with inferiority.
We let $c$ be the sample size in terms of the number of clusters at the interim analysis. The interim decision criteria are defined based on evidence of inferiority with respect to any of the three outcomes. Specifically, the experimental treatment arm will be dropped if sufficient evidence points to inferiority in AE safety, completion, or tolerability. That is, the treatment arm will be dropped if for any $k=1,2,3$,
$\tau_{k}(\mathcal{D}_{c})=\Pr(\delta_{k}(\boldsymbol{\theta})<\delta_{U,k}\mid\mathcal{D}_{c})\leq\gamma_{k}.$
We emphasize that the interim analysis only accommodates stopping based on sufficient evidence of inferiority of the new treatment, which fundamentally differs from the absence of evidence to conclude non-inferiority. Non-inferiority of the new treatment can only be concluded at the final analysis.
We next describe the decision criteria for the final analysis, which will occur once $1.25c$ clusters have been observed. The choice for this constant of 1.25 is based on the motivating trial. If $1.25c\notin\mathbb{Z}^{+}$, the sample size for the final analysis can be rounded up to the nearest integer. At the final analysis, only the posterior probability of non-inferiority related to the AE outcome will be considered: $\Pr(\delta_{1}(\boldsymbol{\theta})<\delta_{U,1}\mid\mathcal{D}_{1.25c})$. We let $\tau_{4}(\mathcal{D}_{c})=\Pr(\delta_{1}(\boldsymbol{\theta})<\delta_{U,1}\mid\mathcal{D}_{1.25c})$ to distinguish between the posterior probability considered at the end of the trial and $\tau_{1}(\mathcal{D}_{c})$ from the interim analysis. The new treatment will be declared non-inferior to the reference arm if $\tau_{4}(\mathcal{D}_{c})>\gamma_{4}$ and the new treatment was continued after the interim analysis. That is, the binary indicator $\nu(\mathcal{D}_{c})$ defined in Section 2.1 equals 1 if and only if $\tau_{k}(\mathcal{D}_{c})>\gamma_{k}$ for all $k=1,\dots,4$.
@@ 5 Application to the SSTaRLeT Trial
We consider four design scenarios for this trial that are summarized in Table 1. The first three scenarios correspond to non-inferiority under $H_{1}$ and the fourth corresponds to inferiority under $H_{0}$. We aim to select a sample size $c$ such that the power to correctly conclude non-inferiority at the final analysis for the acceptable treatment in scenario 2 is at least $\Gamma_{1}=0.8$. The decision thresholds $\boldsymbol{\gamma}$ will be selected to bound the probability of incorrectly concluding non-inferiority for the unacceptable treatment in scenario 4 by $\Gamma_{0}=0.025$. The probabilities of concluding non-inferiority in scenarios 1 (clearly acceptable) and 3 (barely acceptable) are considered for sensitivity analysis.
Table 1: Marginal probabilities for binary outcomes in various design scenarios
@@
We now broadly overview the data that inform the posterior of $\boldsymbol{\delta}(\boldsymbol{\theta})$. We generate a matrix of outcomes ${\bf{Y}}_{N\times 3}$, where each row corresponds to a patient. The $k^{\text{th}}$ column of ${\bf{Y}}_{N\times 3}$ corresponds to the $k^{\text{th}}$ binary outcome: the occurrence of an AE, non-completion, or non-tolerability. There are no additional observed covariates ${\bf{X}}_{N\times p}$ for this example. As such, $\mathcal{D}_{c}=\{{\bf{Y}}_{N\times 3},{\bf{A}}_{N}\}$. Furthermore, we use three intraclass correlation coefficient (ICC) settings – low, moderate, and high – to explore various levels of dependence between the binary outcomes. Since we marginalize over cluster-specific latent effects, the ICC settings materially impact sample size recommendations. It is therefore important to efficiently consider various ICC settings in trial design. We thoroughly describe the data generation procedure for all described settings in Appendix C, and direct readers to Golchi et al. (2022) for additional context on decision criteria based on multiple endpoints.
Each outcome is analyzed using a separate Bayesian logistic regression model with random intercepts:
$y_{k,j,i}\sim\text{BIN}(1,\text{expit}(\eta_{k,j,i})),~{}~{}\text{where}~{}~{}\eta_{k,j,i}=\beta_{0,k}+\beta_{1,k}A_{j,i}+w_{k,j},$ (8)
For the model in (8), the subscripts are such that $k$ indexes the estimand, $j$ indexes the cluster, and $i$ indexes the observation within the $j^{\text{th}}$ cluster. In the model corresponding to the $k^{\text{th}}$ estimand, $\beta_{0,k}$ and $\beta_{1,k}$ represent a common intercept and slope and $w_{k,j}\sim\mathcal{N}(0,\sigma^{2}_{k})$ are cluster-specific intercepts. The latent effects for this example are ${\bf{W}}_{c\times 3}$, where the $k^{\text{th}}$ column contains the random intercepts for the $k^{\text{th}}$ estimand. The parameters $\boldsymbol{\theta}$ consist of the $3\times 1$ vectors of coefficients $\boldsymbol{\beta}_{0}$ and $\boldsymbol{\beta}_{1}$ along with the variances for the random intercepts $\{\sigma^{2}_{k}\}_{k=1}^{3}$. We use diffuse priors for all parameters as described in Appendix C.
Before implementing our approach, we describe how Algorithm 1 can be modified to accommodate this particular design. In Lines 2, 4, and 6 of Algorithm 1, each simulation repetition used to estimate the sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ now has four components: three for the interim analysis and one for the final. For each sample, the data $\mathcal{D}_{1.25c,r}$ are generated by combining $\mathcal{D}_{c,r}$ with data from $0.25c$ new clusters. As in Lines 7 and 8 of Algorithm 1, linear approximations for the final analysis can be constructed based on the logits $\{l_{4}(\mathcal{D}_{c_{0},1,r})\}_{r=1}^{m}$ and $\{l_{4}(\mathcal{D}_{c_{1},1,r})\}_{r=1}^{m}$. However, we must use the final-stage sample sizes $1.25c_{0}$ and $1.25c_{1}$ when obtaining the intercepts and slopes of the linear approximations for the final analysis. The approximation to the joint sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ can be constructed as in Line 9 of Algorithm 1, and no other modifications to our method are required. We direct readers to Hagar et al. (2025) for a discussion of the theory that underlies this extension.
@@
We first present the simulation results for the low ICC setting. When implementing Algorithm 1 with $m=10^{4}$ simulation repetitions, an initial sample size of $c_{0}=120$ was selected based on an anticipated budget for the trial. The probability model $\Psi_{0}$ for this example corresponds to scenario 4 in Table 1. Based on the estimated sampling distribution obtained in Line 2 of Algorithm 1, we chose decision thresholds in Line 3. There are many combinations of decision thresholds that would suitably bound the type I error rate, so we chose thresholds for the interim analysis with a secondary goal of achieving a high probability of stopping the new treatment for inferiority under $\Psi_{0}$. A threshold of $\gamma_{1}=0.8$ corresponding to the primary outcome was first selected. Equal thresholds $\gamma_{2}=\gamma_{3}=0.55$ were then chosen for the secondary outcomes as the smallest thresholds that achieve a 92% probability of correctly stopping for inferiority. Given these interim thresholds, we selected $\gamma_{4}=0.94$ as the smallest threshold that bounded the type I error rate (the probability of an incorrect non-inferiority conclusion) by $\Gamma_{0}=0.025$
@@
The probability model $\Psi_{1}$ for this example corresponds to scenario 2 in Table 1. For this $\Psi_{1}$ model and $c_{0}=120$, the estimated power given $1.25c_{0}$ clusters at the final analysis was 0.6685, which is less than the target probability of $\Gamma_{1}=0.8$. Given that estimate, we next explored a sample size of $c_{1}=180$ to facilitate assessment of the operating characteristics for $c\in[100,200]$. The linear approximations in Lines 7 and 8 were constructed using $c_{0}=120$ and $c_{1}=180$. We note that the probability models $\Psi_{0}$ and $\Psi_{1}$ for this example do not incorporate uncertainty in the parameter values used to generate the data. These choices for $\Psi_{0}$ and $\Psi_{1}$ mirror those used to design SSTaRLeT, but our methodology accommodates nondegenerate $\Psi$ models as described in Section 4. Based on the linear approximations in Algorithm 1, the recommended sample size in Line 10 for the low ICC setting is $c_{2}=162$. A 95% bootstrap confidence interval for the optimal sample size obtained using the procedure detailed in Section 4 with $M=10^{4}$ bootstrap samples is $[150,165]$.
@@
Figure 1: The probability of concluding non-inferiority at the final analysis for the low ICC setting. The black curves are estimated using linear approximations. The dashed curves are pointwise 95% bootstrap confidence intervals. The blue curves arise from simulating sampling distributions for many $c$ values.
@@
Figure 1 visualizes the probability of concluding non-inferiority at the final analysis for each scenario in Table 1 with respect to $c$ for the low ICC setting. This probability corresponds to power for scenarios 1, 2, and 3 and the type I error rate for scenario 4. The black curves were estimated using linear approximations to logits of posterior probabilities at only two sample sizes ($c_{0}=120$ and $c_{1}=180$) using the process in Lines 4 to 9 of Algorithm 1. The dashed curves represent pointwise 95% bootstrap confidence intervals for the design operating characteristics obtained using linear approximations with bootstrap samples as described in Section 4. The blue curves were simulated by independently generating samples of data to estimate sampling distributions of posterior probabilities at $c=\{100,110,\dots,200\}$ for each setting.
@@
Although the blue curves are evidently impacted by simulation variability, we use them as approximate surrogates for the true design operating characteristics. We observe good alignment for all four scenarios between the black curves estimated using linear approximations and the blue ones obtained by independently simulating samples. Furthermore, the blue curves are generally contained within the pointwise 95% bootstrap confidence intervals for the range of sample sizes we consider. However, the black curves are substantially easier to obtain since we need only estimate sampling distributions of posterior probabilities at two values of $c$.
@@
It took roughly 80 minutes on a high-computing server to estimate each black curve in Figure 1 when approximating each posterior using Markov chain Monte Carlo with only one chain of 500 burnin iterations and 2000 retained draws. While initial simulations used multiple Markov chains to verify convergence, our large-scale simulations employed less intensive settings that achieved reasonable convergence. We considered 11 values of $c$ to simulate each blue curve in Figure 1, taking approximately 7 hours using the same computing resources. For standard simulation-based methods, binary search could recommend a sample size $c\in[1,B]$ by estimating the sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ at $\log_{2}(B)$ values of $c$, where $\log_{2}(B)$ is generally much greater than two. Unlike standard methods, our approach also allows practitioners to assess design operating characteristics for new values of $c$ without conducting additional simulations.
To facilitate comparison with the low ICC setting, we repeated the process described above with the same decision thresholds $\boldsymbol{\gamma}=(0.8,0.55,0.55,0.94)$ for the moderate and high ICC settings. Figures 2 and 3 respectively visualize the probabilities of concluding non-inferiority for the moderate and high ICC settings. We again observe good alignment between the black curves estimated using our method and the blue ones obtained by traditional simulation. The computational savings associated with using our methodology are similar for all ICC settings.
@@
Figure 2: The probability of concluding non-inferiority at the final analysis for the moderate ICC setting. The black curves are estimated using linear approximations. The dashed curves are pointwise 95% bootstrap confidence intervals. The blue curves arise from simulating sampling distributions for many $c$ values.
Figure 3: The probability of concluding non-inferiority at the final analysis for the high ICC setting. The black curves are estimated using linear approximations. The dashed curves are pointwise 95% bootstrap confidence intervals. The blue curves arise from simulating sampling distributions for many $c$ values.
@@
Our method recommended a sample size of $c=179$ for the moderate ICC setting; the 95% bootstrap confidence interval for the optimal sample size was [176, 183]. The corresponding sample size recommendation and confidence interval for the high ICC setting were $c=196$ and [192, 202]. These results demonstrate that sample size recommendations are sensitive to the ICC settings when we marginalize over the distributions of the latent effects. Thus, it is crucial to explore various ICC settings during trial design, which increases the number of design configurations considered. Our methodology facilitates the efficient assessment of operating characteristics for each configuration.
@@
To further evaluate the performance of our method, Table 2 summarizes the estimated stopping probabilities at the interim analysis for the recommended sample size corresponding to each ICC scenario. Table 3 presents the estimated power at the final analysis for each scenario. The probabilities in both tables were first estimated using the linear approximations from the extension of Algorithm 1 that were anchored at $c_{0}=120$ and $c_{1}=180$. Each pair of rows compares those probabilities with probabilities obtained by estimating the joint sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ at the recommended sample size $c$. We observe good alignment between each pair of rows, demonstrating the strong performance of our method. In Table 2, we observe that the probability of stopping the unacceptable treatment is roughly 0.92 in all scenarios. Moreover, the stopping probabilities are relatively low for the clearly acceptable and acceptable treatments. Table 3 confirms that the desired power of $\Gamma_{1}=0.8$ and type I error rate of $\Gamma_{0}=0.025$ are reasonably maintained.
@@
Table 2: Estimated stopping probabilities at the interim analysis for the recommended sample sizes. These estimates were obtained using linear approximations from the extension of Algorithm 1 and by simulating confirmatory estimates of sampling distributions.
@@
Low 162 Alg. 1 0.0187 0.1361 0.6204 0.9185
Simulation 0.0194 0.1394 0.6273 0.9200
Moderate 179 Alg. 1 0.0192 0.1409 0.6193 0.9222
Simulation 0.0201 0.1362 0.6175 0.9155
High 196 Alg. 1 0.0199 0.1367 0.6197 0.9157
Simulation 0.0235 0.1462 0.6275 0.9242
Table 3: Estimated power at the final analysis for the recommended sample sizes. These estimates were obtained using linear approximations from the extension of Algorithm 1 and by simulating confirmatory estimates of sampling distributions.
ICC Setting $1.25c$ Method Experimental Treatment
Low 203 Alg. 1 0.9763 0.8004 0.1789 0.0257
Simulation 0.9765 0.7986 0.1754 0.0248
Moderate 224 Alg. 1 0.9773 0.8004 0.1884 0.0228
Simulation 0.9761 0.8064 0.1810 0.0245
High 245 Alg. 1 0.9768 0.8005 0.1790 0.0247
Simulation 0.9730 0.7960 0.1832 0.0253

### v3

@@ 5 Sample Size Determination Procedure
Bootstrap confidence intervals for the design operating characteristics at a given value of $c$ can similarly be constructed. For each of the $M$ sets of bootstrap samples, the linear approximations obtained using the process in Lines 7 to 9 of Algorithm 1 give rise to a new estimate of the operating characteristics. A bootstrap confidence interval for the operating characteristic can also be calculated using the percentile method. In Section 6, we consider the performance of Algorithm 1 and construct bootstrap confidence intervals for our illustrative example.
6 Numerical Studies
We now apply the proposed approach to our illustrative example that is inspired by the SSTARLET trial. In this example, our goal is to show that the new tuberculosis preventative medication is non-inferior to the control regimen with respect to safety. In particular, we aim to support the hypothesis $H_{1}:\delta(\boldsymbol{\theta})<0.04$, where $\delta(\boldsymbol{\theta})$ is the difference between the marginal AE rates in the treatment and control arms. Notation for the data, estimand, and analysis model for this example was introduced in Section 2. Notation for the hypotheses and posterior probabilities was introduced in Section 3.
@@ 6 Numerical Studies
Below, we overview the data-generation processes for this example. We consider four design scenarios for this trial that are summarized in Table 1. The first three scenarios correspond to non-inferiority under $H_{1}$ and the fourth corresponds to inferiority under $H_{0}$. We aim to select a cluster count $c$ such that the power to correctly conclude non-inferiority for the clearly acceptable treatment in scenario 1 is at least $1-\beta=0.8$. The decision threshold $\gamma$ will be selected to bound the type I error rate for the unacceptable treatment in scenario 4 by $\alpha=0.025$. The probabilities of concluding non-inferiority in scenarios 2 (acceptable) and 3 (barely acceptable) are considered for sensitivity analysis.
Table 1: Marginal probabilities for the AE rate in various design scenarios
@@
Furthermore, we use three intraclass correlation coefficient (ICC) settings – low, moderate, and high – to explore various levels of within-cluster dependence between the AE outcomes. Since we marginalize over cluster-specific random effects in this example, the ICC settings materially impact sample size recommendations. It is therefore important to efficiently consider various ICC settings in trial design. In Appendix C, we thoroughly describe the data generation procedure for all described settings and the diffuse analysis priors used for all parameters.
@@
We first present the simulation results for the low ICC setting. When implementing Algorithm 1 with $m=10^{4}$ simulation repetitions, an initial cluster count of $c_{0}=100$ was selected based on an anticipated budget for the trial. The probability model $\Psi_{0}$ for this example corresponds to scenario 4 in Table 1. Based on the estimated sampling distribution obtained in Line 2 of Algorithm 1, we chose $\gamma=0.97$ in Line 3. This threshold was the smallest one rounded to two decimal places that bounded the type I error rate (the probability of an incorrect non-inferiority conclusion) by $\alpha=0.025$
@@
The probability model $\Psi_{1}$ for this example corresponds to scenario 1 in Table 1. For this $\Psi_{1}$ model and $c_{0}=100$, the estimated power was 0.7200, which is less than the target probability of $1-\beta=0.8$. Given that estimate, we next explored a cluster count of $c_{1}=140$ to facilitate assessment of the operating characteristics for $c\in[80,160]$. The linear approximations in Lines 7 to 9 were constructed using $c_{0}=100$ and $c_{1}=140$. We note that the probability models $\Psi_{0}$ and $\Psi_{1}$ for this example do not incorporate uncertainty in the parameter values used to generate the data. These choices for $\Psi_{0}$ and $\Psi_{1}$ mirror those used to design SSTARLET, but our methodology accommodates nondegenerate $\Psi$ models as described in Section 5. Based on the linear approximations in Algorithm 1, the recommended cluster count in Line 11 for the low ICC setting is $c_{2}=115$. A 95% bootstrap confidence interval for the optimal cluster count obtained using the procedure detailed in Section 5 with $M=10^{4}$ bootstrap samples is $[114,116]$.
@@
Figure 1: The probability of concluding non-inferiority for the low ICC setting. The black curves are estimated using linear approximations. The dashed curves are pointwise 95% bootstrap confidence intervals. The blue curves arise from simulating sampling distributions for many $c$ values.
@@
Figure 1 visualizes the probability of concluding non-inferiority at the final analysis for each scenario in Table 1 with respect to $c$ for the low ICC setting. This probability corresponds to power for scenarios 1, 2, and 3 and the type I error rate for scenario 4. The black curves were estimated using linear approximations to logits of posterior probabilities at only two cluster counts ($c_{0}=100$ and $c_{1}=140$) using the process in Lines 4 to 9 of Algorithm 1. The dashed curves represent pointwise 95% bootstrap confidence intervals for the design operating characteristics obtained using linear approximations with bootstrap samples as described in Section 5. The blue curves were simulated by generating samples of data to estimate sampling distributions of posterior probabilities at $c=\{80,90,\dots,160\}$ for each setting.
@@
Although the blue curves are evidently impacted by simulation variability, we use them as approximate surrogates for the true design operating characteristics. We observe good alignment for all four scenarios between the black curves estimated using linear approximations and the blue ones obtained by separately simulating samples. Furthermore, the blue curves are generally contained within the pointwise 95% bootstrap confidence intervals for the range of cluster counts we consider. However, the black curves are substantially easier to obtain since we need only estimate sampling distributions of posterior probabilities at two values of $c$.
@@
It took roughly 8 minutes on a high-computing server to estimate each black curve in Figure 1 when approximating each posterior using Markov chain Monte Carlo with only one chain of 500 burnin iterations and 2000 retained draws. While initial simulations used multiple Markov chains to verify convergence, our large-scale simulations employed less intensive settings that achieved reasonable convergence. We considered 9 values of $c$ to simulate each blue curve in Figure 1, taking approximately 35 minutes using the same computing resources. For standard simulation-based methods, binary search could recommend a cluster count $c\in[1,B]$ by estimating the sampling distribution of $\tau(\mathcal{D}_{c})$ at $\log_{2}(B)$ values of $c$, where $\log_{2}(B)$ is generally much greater than two. Unlike standard methods, our approach also allows practitioners to assess design operating characteristics for new values of $c$ without conducting additional simulations.
@@
Figure 2: The probability of concluding non-inferiority for the moderate ICC setting. The black curves are estimated using linear approximations. The dashed curves are pointwise 95% bootstrap confidence intervals. The blue curves arise from simulating sampling distributions for many $c$ values.
To facilitate comparison with the low ICC setting, we repeated the process described above with the same decision threshold $\gamma=0.97$ for the moderate and high ICC settings. Figures 2 and 3 respectively visualize the probability of concluding non-inferiority for the moderate and high ICC settings. We again observe good alignment between the black curves estimated using our method and the blue ones obtained by traditional simulation. The computational savings associated with using our methodology are similar for all ICC settings.
@@
Our method recommended a cluster count of $c=119$ for the moderate ICC setting; the 95% bootstrap confidence interval for the optimal cluster count was [118, 120]. The corresponding cluster count recommendation and confidence interval for the high ICC setting were $c=129$ and [128, 131]. These results demonstrate that cluster count recommendations are sensitive to the ICC settings when we marginalize over the distributions of the random effects. Thus, it is crucial to explore various ICC settings during trial design, which increases the number of design configurations considered. Our methodology facilitates the efficient assessment of operating characteristics for each configuration.
Figure 3: The probability of concluding non-inferiority for the high ICC setting. The black curves are estimated using linear approximations. The dashed curves are pointwise 95% bootstrap confidence intervals. The blue curves arise from simulating sampling distributions for many $c$ values.
@@
To further evaluate the performance of our method, Table 2 summarizes the estimated probabilities of concluding non-inferiority for the recommended cluster count corresponding to each ICC scenario. The probabilities were first estimated using the linear approximations from Algorithm 1 that were anchored at $c_{0}=100$ and $c_{1}=140$. Each pair of rows compares those probabilities with probabilities obtained by estimating the sampling distribution of $\tau(\mathcal{D}_{c})$ at the recommended cluster count $c$. We observe good alignment between each pair of rows, demonstrating the strong performance of our method. Table 2 confirms that the desired power of $1-\beta=0.8$ and type I error rate of $\alpha=0.025$ are reasonably maintained.
@@
Table 2: Estimated probabilities of concluding non-inferiority at the recommended cluster counts. These estimates were obtained using linear approximations from Algorithm 1 and by simulating confirmatory estimates of sampling distributions.
@@
Low 115 Alg. 1 0.8020 0.5025 0.0793 0.0235
Simulation 0.8110 0.5118 0.0828 0.0212
Moderate 119 Alg. 1 0.8015 0.4923 0.0842 0.0257
Simulation 0.8036 0.4911 0.0811 0.0249
High 129 Alg. 1 0.8016 0.4974 0.0852 0.0279
Simulation 0.8078 0.4967 0.0803 0.0286
We recognize that many cluster-randomized trials do not have as many clusters as our illustrative example based on SSTARLET. In Appendix D of the supplement, we adjust several of the settings for this example to consider the performance of our method with smaller cluster counts. The numerical studies in Appendix D confirm the suitable performance of our method in such settings.
7 Extensions for More Complex Designs
The economical design methodology proposed in this paper for trials with clustered data can be extended to accommodate more complex designs. Several such extensions are needed for the design of the actual SSTARLET trial. For example, the design framework of Hagar and Stevens (2025), which we build upon here, has previously been extended in Hagar et al. (2025) to group sequential designs involving a single estimand and independent observations. In such designs, an alternative hypothesis $H_{1}$ similar to that in (4) is considered across $T$ potential stages of the trial. Criteria for the operating characteristics of group sequential designs, including power and the family-wise error rate, are defined with respect to the joint sampling of posterior summaries across these $T$ analyses. Particularly for adaptive trials that allow early stopping for both efficacy and futility, the operating characteristics require more complex definitions than those we considered in (6) and (7).
To accommodate sequential decision criteria, Hagar et al. (2025) developed theory to account for dependence in the joint sampling distribution of posterior and posterior predictive probabilities across trial stages. That theory was substantiated using a proxy sampling distribution based on the BvM theorem and the joint sampling distribution of the MLE arising from the joint canonical distribution (Jennison and Turnbull, 1999). This joint sampling distribution of the MLE is a $T$-dimensional analog to the $\mathcal{N}(\delta_{r},c^{-1}\Lambda(\boldsymbol{\theta}_{r}))$ distribution that we considered in Section 4. Combining the results from Hagar et al. (2025) with our results here provides an efficient approach to design adaptive cluster trials with a single estimand, where the joint sampling distribution of posterior probabilities is estimated at only two values of $c$. In such an approach, SSD is conducted in terms of the cluster count under the constraint that the proportion of new participants in each trial stage is held fixed, and decision criteria are optimized using the sampling distribution estimates.
More recently, Hagar and Maleyeff et al. (2025) extended the results from Hagar et al. (2025) to design platform trials with multiple estimands in non-clustered settings. That work also accounts for (i) fixed delays between reaching interim recruitment thresholds and performing the corresponding analyses and (ii) response-adaptive randomization resulting from dropping treatment arms. Each of the $K>1$ estimands in the trial is associated with a set of hypotheses, giving rise to multiple complementary hypotheses $\{H_{0,k}\}_{k=1}^{K}$ and $\{H_{1,k}\}_{k=1}^{K}$. Multiple hypotheses further complicate the definition of decision criteria and design operating characteristics. Suppose that all $K$ estimands are considered in each of the $T$ potential stages of the trial. The operating characteristics for platform trials are defined with respect to the joint sampling distribution of posterior probabilities across these $K$ estimands and $T$ analyses.
One contribution of Hagar and Maleyeff et al. (2025) involved deriving new theory to account for dependence in the joint sampling distribution of posterior probabilities across estimands. Their theory leveraged the BvM theorem and the joint sampling distribution of the MLE $\hat{\boldsymbol{\delta}}\in\mathbb{R}^{KT}$. Due to the fixed delays mentioned above, the covariance matrix of this sampling distribution does not take the form $c^{-1}\boldsymbol{\Lambda}(\boldsymbol{\theta}_{r})$, where the cluster count (or sample size for independent data) factors out of the denominator. This factorization was required to create our proxy sampling distribution of posterior probabilities in Section 4, so more intricate theory is needed to construct proxy sampling distributions for platform trials and explore linear trends. Nevertheless, the theory in our paper can be combined with a simplified version of the SSD methodology from Hagar and Maleyeff et al. (2025) to design non-platform cluster trials with multiple endpoints.